# Enhanced Ground–Satellite Direct Access via Onboard Rydberg Atomic Quantum Receivers

### v1

@@ V-B RAQR for Integrated Sensing and Communications
The inherent frequency agility and sensitivity of RAQRs also open the door to integrated sensing and communication (ISAC) in spaceborne platforms. A key direction is using RAQRs onboard small satellites to simultaneously support uplink reception and remote sensing based on reflected or scattered signals of opportunity. Distributed RAQR-equipped satellites may collaboratively decode weak signals from Earth terminals or measure environmental parameters without dedicated radar payloads. Their SI-traceability and polarization-resolving capability further enable precise interference analysis and electromagnetic situational awareness. A promising use case is joint radar-communication operation, where a RAQR senses echoes of communication signals for passive imaging while maintaining link performance. Designing such dual-use protocols and addressing waveform compatibility and sensitivity trade-offs will be essential.
V-C System-Level Challenges and Deployment Pathways
Successful orbital deployment of RAQRs requires coordinated advances in hardware integration, space qualification, and mission validation. Research should focus on photonic and microfabrication techniques that shrink chip‑scale lasers, vapor cells, and optics to CubeSat‑class SWaP. Designers must also solve thermal control, radiation tolerance, and structural stability in the space environment. Operational readiness will depend on robust Doppler‑shift calibration, remote frequency tuning, and seamless insertion of RAQR front ends into standard satellite protocols. A staged flight‑test program will begin with high‑altitude balloons and progress to orbital demonstrators to verify link budgets and end‑to‑end performance.
@@ VI Conclusion
This study demonstrated the feasibility of RAQRs as satellite front ends for 6G space-ground direct access. A review of their quantum‑electrometry principles confirmed their compatibility with space platforms. We proposed an atomic‑electronic hybrid transceiver and examined both narrowband and wideband modes. System‑level simulations reported gains in data rate, coverage, and sensing accuracy over conventional RF receivers, and a scalable architecture extended these benefits to distributed constellations. Key research priorities now include advanced signal processing, ISAC, and space‑qualified hardware that will move RAQR technology from laboratory demonstrations to operational payloads.

### v2

@@ V-B RAQR for Integrated Sensing and Communications
The inherent frequency agility and sensitivity of RAQRs also open the door to integrated sensing and communication (ISAC) on spaceborne platforms, yet they also pose significant self-interference challenges. A key direction is using RAQRs onboard small satellites to simultaneously support uplink reception and remote sensing based on frequency-division duplexing. Distributed RAQR-equipped satellites may collaboratively decode weak signals from Earth terminals or measure environmental parameters without dedicated radar payloads. Their SI-traceability and polarization-resolving capability further enable precise interference analysis and electromagnetic situational awareness. A promising use case is joint radar-communication operation, where a RAQR senses echoes of communication signals for passive imaging while maintaining link performance. Designing such dual-use protocols and addressing waveform compatibility and sensitivity trade-offs will be essential.
V-C Practical Challenges and Deployment Pathways
Successful orbital deployment of RAQRs will require coordinated advances in hardware integration, space qualification, and mission validation. The research should address CubeSat-level SWaP constraints by advancing photonic and microfabrication technologies for miniaturized chip-scale lasers, vapor cells, and optical components. Critical engineering challenges include thermal management, radiation hardening, and structural stability under the space environment. Moreover, Doppler‑shift calibration, remote frequency tuning, and seamless insertion of RAQR front ends into standard satellite protocols need to be tackled in the future. A staged flight‑test program will begin with high‑altitude balloons and progress to orbital demonstrators to verify link budgets and end‑to‑end performance.
@@ VI Conclusion
This study demonstrated the feasibility of RAQRs as satellite front ends for 6G space-ground direct access. A review of their quantum‑electrometry principles confirmed their compatibility with space platforms. We proposed an atomic‑electronic hybrid transceiver and examined both narrowband and wideband modes. System‑level simulations reported gains in data rate, coverage, and sensing accuracy over conventional RF receivers, and a scalable architecture extended these benefits to distributed constellations. Key challenges are outlined, including advanced signal processing, ISAC, and hardware implementation, which collectively pave the way for moving RAQR technology from lab demos to flight-ready payloads.
Acknowledgment
The work of Pei Xiao, Qihao Peng, and Qu Luo was supported in part by the U.K. Engineering and Physical Sciences Research Council under Grant EP/X013162/1. The work of Chau Yuen and Tierui Gong was supported by the Ministry of Education (MoE), Singapore, under its MoE AcRF Tier 1 Thematic Grant RT12/23 023780-00001.